# Efficient Feature Aggregation and Scale-Aware Regression for Monocular 3D Object Detection

### v2

@@ Abstract
Index Terms:
3D object detection, monocular, scale-aware, scene understanding, autonomous driving.
@@ I Introduction
In recent years, 3D object detection has emerged as a pivotal area of research, particularly driven by its critical role in autonomous systems, robotics, and augmented reality. The ability to accurately detect and locate objects in three-dimensional space is essential for applications that require a high degree of spatial awareness, especially for autonomous driving systems[1],[2].
@@
Advancements in sensors such as LiDAR, radar, and stereo cameras have enhanced 3D object detection with point cloud data, radar signals, and depth maps. However, LiDAR’s high-resolution point clouds[3],[4],[5],[6] and stereo methods’ depth maps[7],[8] require additional calibration. CAD[9],[10] and multi-frame[11] approaches improve shape and movement tracking, but are also computationally intensive. Compared to the above methods, monocular vision systems offer a more cost-effective and deployable solution. However, they cannot accurately recover 3D information from 2D images without explicit depth perception.
@@
Therefore, current advances in monocular 3D detection have focused on improving depth estimation accuracy. Methods like[12],[13],[14],[15],[16],[17] following the CenterNet[18] paradigm, use direct regression to estimate the 3D center depth of the target object.[9],[19],[20],[21],[22] incorporate geometric constraints into depth estimation to enhance target depth recovery.[23],[24],[25],[26],[27] effectively integrate geometric depth with regressed depth, improving the network’s generalization ability. However, these approaches focus on decoupling the regression of 2D and 3D attributes: predicting 2D bounding box parameters $\left({x,y,w,h}\right)$ and 3D object parameters $\left({x,y,z,w,h,l,yaw}\right)$ separately. This separation may neglect valuable spatial cues provided by the 2D predictions, such as the object’s size and position in the image, which can serve as geometric information for estimating depth and 3D structure. For instance, distant objects tend to produce smaller 2D bounding boxes, while nearby objects appear larger, indicating a potential correlation between 2D scale and 3D distance that is not fully exploited in decoupled frameworks. This issue is evident in the contrast between the attention heatmaps of the networks shown in Fig. 1. It can be seen that DEVIANT[22] and MonoLSS[17] struggle to focus on smaller, distant objects, particularly in the pedestrian category. Moreover, DEVIANT[22] tends to inadvertently take care of irrelevant background noise. Some previous works attempted to address this issue. Chen et al. proposed a shape-aware auxiliary training task[28]. [29] integrated deformable convolutions[30] to enhance the adaptability of the model to the features. Although these methods achieve dynamic receptive field adjustment to some extent, they have not explicitly considered object scale within a scene, nor dynamically adjusted the model’s attention based on different scales.
@@
Figure 1: Visualization of attention heatmaps for different models: (a) DEVIANT[22], (b) MonoLSS[17], and (c) Our proposed MonoASRH. Previous methods struggle to capture distant and occluded objects due to fixed receptive fields. In contrast, MonoASRH dynamically adjusts attention across multiple scales, improving detection of various object categories, such as cars and pedestrians.
To address the aforementioned limitations, this paper proposes the novel Efficient Hybrid Feature Aggregation Module (EH-FAM) and Adaptive Scale-Aware 3D Regression Head (ASRH). Inspired by RT-DETR’s[31] success in 2D detection, EH-FAM integrates Vision Transformers with CNNs to efficiently aggregate visual features, significantly reducing model complexity compared to traditional methods like DLAUp[32]. The EH-FAM employs self-attention mechanisms for within-scale feature interaction at the highest level, while effectively propagating low-level feature information across all scales through a dedicated CNN architecture. Subsequently, ASRH encodes 2D bounding box dimensions to capture scale features, which are fused with semantic features extracted from the cross-scale feature through a specially designed scale-semantic feature fusion module. The fused features guide the learning of receptive field offsets, which are then applied to deformable convolutions, enhancing the model’s ability to adapt to varying object scales. Furthermore, ASRH utilizes a spatial variance-based attention mechanism to adaptively differentiate foreground objects from noise within semantic features. Additionally, we introduce the Selective Confidence-Guided Heatmap Loss to facilitate ASRH to prioritize high confidence detections and mitigate the impact of hard samples.
@@
  • Extensive experiments on the KITTI 3D object detection benchmark[33] and Waymo Open dataset[34] show that our MonoASRH achieves better results compared to previous state-of-the-art methods.
@@ II-A Monocular 3D Object Detection
In general, Monocular 3D object detection methods can be categorized into three main approaches: depth map-based, center-based, and transformer-based methods. Depth map-based approaches address spatial ambiguity by estimating depth maps using supervised [35],[36] or self-supervised networks[37], [38].[39], with techniques like pseudo-LiDAR conversions[40],[41] and multi-scale depth fusion[42],[43] improving geometric reasoning. However, their reliance on pre-trained depth networks and overlap between training and validation datasets[44] often lead to foreground depth inaccuracies.
Center-based methods extend anchor-free frameworks like CenterNet [18], linking 3D attributes to object centers. To mitigate depth errors, works like[14],[20],[22] integrate geometric constraints and hierarchical learning, while[24] leverages multiple depth attributes to generate various depth estimates. [28],[45] propose shape-aware schemes and plane-constrained 3D detection frameworks, respectively. MonoCD[27] introduces the concept of complementary depth. MonoLSS[17] uses Gumbel-Softmax probabilistic sampling[46] to distinguish between positive and negative samples in depth maps. Despite efficiency, these methods risk overlooking 2D-3D geometric consistency.
Transformer-based models leverage end-to-end architectures[47],[5] for global feature interaction, injecting global depth information into the Transformer to guide detection. MonoDETR[48] interacts with depth and visual features through a depth cross-attention layer in the decoder. To improve inference efficiency, MonoATT[49] introduces an adaptive token Transformer. However, their computational complexity, sensitivity to background noise, and slow convergence hinder real-time applicability.
In this work, we propose a center-based framework that integrates lightweight convolutional modules with multi-head attention for feature aggregation. To enhance 3D detection robustness, we introduce a latent space projection paradigm where 2D scale features are mapped to guide the 3D regression heads.
@@ II-B Feature Aggregation Strategy
Most center-based monocular 3D detection frameworks[14],[20],[22],[27],[17] employ DLAUp[32] as the feature aggregation module. DLAUp’s hierarchical aggregation through pure convolution operations lacks sufficient global perception, and its large-kernel transposed convolution for upsampling also increases network parameters. Although MonoDETR[48] introduces an end-to-end Transformer architecture, traditional Transformer Encoder’s sensitivity to background noise and slow convergence hinder its performance.
Recent studies focus on hybrid CNN-Transformer architectures for feature aggregation. Works like U-shaped hybrid network by[50], cross-scale token attention by[51], dual-resolution self-attention by[52], and RT-DETR’s[31] efficient hybrid encoder demonstrate their effectiveness in modeling global dependencies while preserving local feature extraction. This hybrid architecture can effectively mitigate the aforementioned challenges in feature aggregation modules of monocular 3D detection methods.
Inspired by RT-DETR, our EH-FAM employs multi-head attention on highest-level features for long-range dependency capture. We further design a lightweight CNN-based cross-scale fusion structure that significantly enhances inference efficiency through re-parameterization[53] and convolutional decoupling[54] techniques.
@@ II-C Scale-Aware Object Detection
Scale-aware techniques enhance detection models by incorporating scale information. SAFR-CNN[55] introduces scale-aware weighting to prioritize sub-networks specialized for input scales. SFDet[56] adapts this concept to face detection by associating multi-scale layers with appropriately sized anchors. Chen et al.[57] develop Scale-aware AutoAug for learning data augmentation strategies, while AdaZoom[58] proposes adaptive region zooming based on object scales. SMT[59] advances this field with Scale-Aware Modulation, integrating multi-head hybrid convolutions with lightweight aggregation for context-aware token refinement.
Currently, most methods focus on scale-awareness in 2D object detection, whereas there is relatively little work on monocular 3D detection. In fact, scale information is crucial for 3D detectors, since an object’s 2D size often has an implicit geometric projection relationship with its 3D position. Our ASRH module encodes 2D scale cues to generate dynamic receptive field offsets that guide the regression head toward scale-aware 3D localization.
@@ III Proposed Methodology
In this section, we introduce our MonoASRH architecture. The overall framework, shown in Fig. 2, mainly consists of backbone, Efficient Hybrid Feature Aggregation Module, 2D regression head, and Adaptive Scale-Aware 3D Regression Head. Our pipeline mainly builds on the GUPNet [21]. Detailed implementation will be discussed later in this section.
@@ III-A Overall Architecture
For a given RGB image $I\in{\mathbb{R}^{H\times W\times 3}}$, we extract multi-scale deep features using a pre-trained DLA-34 backbone[32]. These features, represented as $\left\{{{S_{1}},{S_{2}},{S_{3}},{S_{4}}}\right\}$, undergo further refinement and cross-scale fusion through the proposed EH-FAM. The resulting aggregated features are denoted as $F\in{\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times C}}$. Similar to[21], our 2D detector is based on CenterNet[18]. We feed the deep features $F$ into three 2D detection heads to regress heatmaps $H\in{\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times 3}}$, 2D offsets ${O_{2D}}\in{\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times 2}}$, and 2D sizes ${S_{2D}}\in{\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times 2}}$, enabling the prediction of 2D object centers and dimensions.
@@ III-B Efficient Hybrid Feature Aggregation Module
${{Q_{i}}=\left({{X_{seq}}+{P_{e}}}\right)W_{i}^{\left(Q\right)},}&{{K_{i}}=\left({{X_{seq}}+{P_{e}}}\right)W_{i}^{\left(K\right)},}\end{array}$ (1)
@@
2) Upsampling Block: To improve EH-FAM’s ability to perceive small-scale objects, we design an upsample block and fusion block composed of lightweight convolutional modules. These blocks enable cross-scale information interaction and enhance the extraction of local semantic features. In the upsample block, we replace DLAUp’s[32] large-kernel transposed convolutions with bilinear upsampling. To further mitigate information loss, we apply a 7×7 convolution prior to upsampling, which effectively refines feature maps. For efficiency, we decouple the 7$\times$7 convolution into successive 1$\times$7 and 7$\times$1 convolutions, following the approach in[54]. Additionally, since some categories, such as Pedestrian and Cyclist, have relatively large aspect ratios but small sizes, decoupled convolutions in the horizontal and vertical directions are adopted to capture these elongated scale features.
3) Fusion Block: In the fusion block, we employ the RepVGGplus block[53], which leverages re-parameterization techniques to convert a multi-branch ResNet-style architecture into a single-path VGG-like model during inference. This maintains the representational power of the model while accelerating inference. The final deep features are aggregated as follows:
@@
where ${f}$ denotes the upsampling and cross-scale fusion operations.
@@ III-C Adaptive Scale-Aware 3D Regression Head
2) Scale-Semantic Feature Fusion Module: This module comprises two key components: offset feature generation and attention mask generation. For offset generation, the scale feature ${G_{o}}$ and local semantic features ${F_{R}}$ are concatenated. The combined features are processed through a stacked convolutional layer consisting of 3$\times$3 and 1$\times$1 convolutions, producing a scale-aware offset ${\Delta_{S}}$, which is used to dynamically adjust the receptive field in subsequent deformable convolutions.
@@
In the attention mask generation component, inspired by[60], we design an attention mechanism based on spatial variance. This mechanism emphasizes regions with significant deviations from the mean, which are often foreground regions of interests. First, the mean of the local semantic features ${F_{R}}$ are computed across spatial dimensions:
@@
3) 3D Regression Head: To achieve scale-aware dynamic receptive field adjustment, the first layer of the 3D regression head utilizes deformable convolutions[30]. As shown in Eqn. 13, the offset ${\Delta_{S}}$ and attention mask $M$ generated by the Scale-Semantic Feature Fusion Module are applied to this layer, allowing the model to better detect and localize objects of varying scales:
@@
where $\Omega$ represents the local neighborhood for a 3$\times$3 kernel, ${{P_{n}}}$ enumerates the locations in $\Omega$, ${\Delta_{S}^{{P_{n}}}}$ is the learned offset for each position in the convolution kernel, ${{W_{{P_{n}}}}}$ is the weight of the convolution kernel at position ${{P_{n}}}$, and ${{M_{{P_{n}}}}}$ controls the contribution of the corresponding kernel position.
Next, we normalize the feature map ${F_{de}}$ using the Attentive Normalization layer[61], which is a light weight module integrating feature normalization and channel-wise feature attention:
@@
Due to AN’s mixture modeling approach to affine transformation, which recalibrates features after standardization, it is employed in the regression head to facilitate the learning of more expressive latent feature representations. Finally, the feature map is processed through a LeakyReLU activation function and a 1$\times$1 convolution for channel mapping, ultimately regressing the 3D bounding box attributes.

### v3

@@ I Introduction
In recent years, 3D object detection has emerged as a pivotal area of research, particularly driven by its critical role in autonomous systems, robotics, and augmented reality. The ability to accurately detect and locate objects in three-dimensional space is essential for applications that require a high degree of spatial awareness, especially for autonomous driving systems[arnold2019survey],[grigorescu2020survey].
@@
Advancements in sensors such as LiDAR, radar, and stereo cameras have enhanced 3D object detection with point cloud data, radar signals, and depth maps. However, LiDAR’s high-resolution point clouds[chen2021monorun],[reading2021categorical],[huang2022monodtr],[chong2022monodistill] and stereo methods’ depth maps[li2020confidence],[zhou2022sgm3d] require additional calibration. CAD[liu2021autoshape],[li2022densely] and multi-frame[brazil2020kinematic] approaches improve shape and movement tracking, but are also computationally intensive. Compared to the above methods, monocular vision systems offer a more cost-effective and deployable solution. However, they cannot accurately recover 3D information from 2D images without explicit depth perception.
Figure 1: Schematic representation of vehicle imaging size at varying distances from the camera. Cars farther from the camera tend to occupy a smaller proportion of the image, while those closer tend to occupy a larger proportion.
@@
Therefore, current advances in monocular 3D detection have focused on improving depth estimation accuracy. Methods like[zhou2019objects],[liu2020smoke],[ma2021delving],[luo2021m3dssd],[liu2022learning],[li2024monolss] following the CenterNet[duan2019centernet] paradigm, use direct regression to estimate the 3D center depth of the target object.[liu2021autoshape],[li2020rtm3d],[zhang2021objects],[lu2021geometry],[kumar2022deviant] incorporate geometric constraints into depth estimation to enhance target depth recovery.[kim2022boosting],[li2022diversity],[qin2022monoground],[zhu2023monoedge],[yan2024monocd] effectively integrate geometric depth with regressed depth, improving the network’s generalization ability. However, these approaches focus on decoupling the regression of 2D and 3D attributes: predicting 2D bounding box parameters $\left\{{{x_{2D}},{y_{2D}},{w_{2D}},{h_{2D}}}\right\}$ and 3D object parameters $\left\{{{x_{3D}},{y_{3D}},{z_{3D}},{w_{3D}},{h_{3D}},{l_{3D}},yaw}\right\}$ separately. This separation may neglect valuable spatial cues provided by the 2D predictions, such as the object’s size and position in the image, which can serve as geometric information for estimating depth and 3D structure. For instance, distant objects tend to produce smaller 2D bounding boxes, while nearby objects appear larger (Fig. 1), indicating a potential correlation between 2D scale and 3D distance that is not fully exploited in decoupled frameworks. This issue is evident in the contrast between the attention heatmaps of the networks shown in Fig. 2. It can be seen that DEVIANT[kumar2022deviant] and MonoLSS[li2024monolss] struggle to focus on smaller, distant objects, particularly in the pedestrian category. Moreover, DEVIANT[kumar2022deviant] tends to inadvertently take care of irrelevant background noise. Some previous works attempted to address this issue. Chen et al. proposed a shape-aware auxiliary training task[chen2023shape]. [chen2023monocular] integrated deformable convolutions[zhu2019deformable] to enhance the adaptability of the model to the features. Although these methods achieve dynamic receptive field adjustment to some extent, they have not explicitly considered object scale within a scene, nor dynamically adjusted the model’s attention based on different scales.
@@
Figure 2: Visualization of attention heatmaps for different models: (a) DEVIANT[kumar2022deviant], (b) MonoLSS[li2024monolss], and (c) Our proposed MonoASRH. Previous methods struggle to capture distant and occluded objects due to fixed receptive fields. In contrast, MonoASRH dynamically adjusts attention across multiple scales, improving detection of various object categories, such as cars and pedestrians.
To address the aforementioned limitations, this paper proposes the novel Efficient Hybrid Feature Aggregation Module (EH-FAM) and Adaptive Scale-Aware 3D Regression Head (ASRH). Inspired by RT-DETR’s[zhao2024detrs] success in 2D detection, EH-FAM integrates Vision Transformers with CNNs to efficiently aggregate visual features, significantly reducing model complexity compared to traditional methods like DLAUp[yu2018deep]. The EH-FAM employs self-attention mechanisms for within-scale feature interaction at the highest level, while effectively propagating low-level feature information across all scales through a dedicated CNN architecture. Subsequently, ASRH encodes 2D bounding box dimensions to capture scale features, which are fused with semantic features extracted from the cross-scale feature through a specially designed scale-semantic feature fusion module. The fused features guide the learning of receptive field offsets, which are then applied to deformable convolutions, enhancing the model’s ability to adapt to varying object scales. Furthermore, ASRH utilizes a spatial variance-based attention mechanism to adaptively differentiate foreground objects from noise within semantic features. Additionally, we introduce the Selective Confidence-Guided Heatmap Loss to facilitate ASRH to prioritize high confidence detections and mitigate the impact of hard samples.
@@
  • Extensive experiments on the KITTI 3D object detection benchmark[geiger2012we] and Waymo Open dataset[sun2020scalability] show that our MonoASRH achieves better results compared to previous state-of-the-art methods.
@@ II-A Monocular 3D Object Detection
In general, Monocular 3D object detection methods can be categorized into three main approaches: depth map-based, center-based, and transformer-based methods. Depth map-based approaches address spatial ambiguity by estimating depth maps using supervised [fu2018deep],[lee2019big] or self-supervised networks[godard2019digging], [guizilini20203d].[ding2020learning], with techniques like pseudo-LiDAR conversions[weng2019monocular],[ma2019accurate] and multi-scale depth fusion[wang2021depth],[ma2020rethinking] improving geometric reasoning. However, their reliance on pre-trained depth networks and overlap between training and validation datasets[simonelli2021we] often lead to foreground depth inaccuracies.
Center-based methods extend anchor-free frameworks like CenterNet [duan2019centernet], linking 3D attributes to object centers. To mitigate depth errors, works like[ma2021delving],[zhang2021objects],[kumar2022deviant] integrate geometric constraints and hierarchical learning, while[li2022diversity] leverages multiple depth attributes to generate various depth estimates. [chen2023shape],[yao2023occlusion] propose shape-aware schemes and plane-constrained 3D detection frameworks, respectively. MonoCD[yan2024monocd] introduces the concept of complementary depth. MonoLSS[li2024monolss] uses Gumbel-Softmax probabilistic sampling[jang2016categorical] to distinguish between positive and negative samples in depth maps. Despite efficiency, these methods risk overlooking 2D-3D geometric consistency.
Transformer-based models leverage end-to-end architectures[wang2022detr3d],[huang2022monodtr] for global feature interaction, injecting global depth information into the Transformer to guide detection. MonoDETR[zhang2023monodetr] interacts with depth and visual features through a depth cross-attention layer in the decoder. To improve inference efficiency, MonoATT[zhou2023monoatt] introduces an adaptive token Transformer. However, their computational complexity, sensitivity to background noise, and slow convergence hinder real-time applicability.
In this work, we propose a center-based framework called MonoASRH that integrates lightweight convolutional modules with multi-head attention for feature aggregation. To further enhance 3D detection robustness, we introduce a latent space projection paradigm where 2D scale features are mapped to guide the 3D regression heads.
@@ II-B Feature Aggregation Strategy
Predominant center-based monocular 3D detection frameworks[ma2021delving],[zhang2021objects],[kumar2022deviant],[yan2024monocd],[li2024monolss] utilize DLAUp[yu2018deep] for feature aggregation. However, DLAUp’s hierarchical architecture relies solely on convolutional operations, limiting its capacity for global context integration. Additionally, its dependence on large-kernel transposed convolutions for upsampling significantly increases parametric complexity. Although MonoDETR[zhang2023monodetr] introduces an end-to-end Transformer architecture to address these limitations, the conventional Transformer Encoder struggles with sensitivity to irrelevant background features and exhibits slow convergence, ultimately constraining detection performance.
Recent research has increasingly focused on hybrid CNN-Transformer architectures for feature aggregation. Approaches such as the U-shaped hybrid network[deng2023cformer], cross-scale token attention mechanisms[yoo2023enriched], dual-resolution self-attention designs[ilyas2024hybrid], and RT-DETR’s efficient hybrid encoder[zhao2024detrs] collectively demonstrate the capability of such architectures to model long-range dependencies while retaining localized feature extraction. This hybrid architecture can effectively mitigate the aforementioned challenges in feature aggregation modules of monocular 3D detection methods.
Inspired by RT-DETR, our proposed EH-FAM architecture incorporates a multi-head attention mechanism to process highest-level feature maps. This design captures long-range dependencies within abstract semantic representations, enabling comprehensive global context modeling. Furthermore, we design a lightweight CNN-based cross-scale feature fusion structure that significantly enhances inference efficiency through re-parameterization[ding2021repvgg] and convolutional decoupling[szegedy2017inception] techniques.
@@ II-C Scale-Aware Object Detection
Scale-aware methodologies improve detection robustness by explicitly incorporating dimensional information. SAFR-CNN[li2017scale] introduces scale-aware weighting to prioritize sub-networks specialized for input scales. Extending this paradigm, SFDet[zhang2019single] adapts the strategy for face detection through systematic association of multi-scale feature layers. Complementary research by Chen et al.[chen2021scale] develops Scale-aware method to optimize augmentation policies, whereas AdaZoom[xu2022adazoom] introduces adaptive region zooming conditioned on object scale characteristics. SMT[lin2023scale] further advances this field with Scale-Aware Modulation, integrating hybrid convolutions with lightweight aggregation for context-aware token refinement.
Currently, most methods focus on scale awareness in 2D object detection, whereas there is relatively little work on monocular 3D detection. Notably, scale information constitutes a critical geometric prior for 3D detectors due to the inherent relationship between an object’s projected 2D dimensions and its spatial position in 3D space. To address this limitation, our ASRH module encodes 2D scale cues to generate dynamic receptive field offsets that guide the regression head toward scale-aware 3D localization.
@@ III Proposed Methodology
In this section, we introduce our MonoASRH architecture. The overall framework, shown in Fig. 3, mainly consists of backbone, Efficient Hybrid Feature Aggregation Module, 2D regression head, and Adaptive Scale-Aware 3D Regression Head. Our pipeline mainly builds on the GUPNet [lu2021geometry]. Detailed implementation will be discussed later in this section.
@@ III-A Overall Architecture
For a given RGB image $I\in{\mathbb{R}^{H\times W\times 3}}$, we extract multi-scale deep features using a pre-trained DLA-34 backbone[yu2018deep]. These features, represented as $\left\{{{S_{1}},{S_{2}},{S_{3}},{S_{4}}}\right\}$, undergo further refinement and cross-scale fusion through the proposed EH-FAM. The resulting aggregated features are denoted as $F\in{\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times C}}$. Similar to[lu2021geometry], our 2D detector is based on CenterNet[duan2019centernet]. We feed the deep features $F$ into three 2D detection heads to regress heatmaps $H\in{\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times 3}}$, 2D offsets ${O_{2D}}\in{\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times 2}}$, and 2D sizes ${S_{2D}}\in{\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times 2}}$, enabling the prediction of 2D object centers and dimensions.
@@ III-B Efficient Hybrid Feature Aggregation Module
[…]
@@
[…]
Figure 5: Structural diagram of the RepVGGplus block. During training, RepVGGplus employs a multi-branch convolutional architecture, which is re-parameterized into a single $\times3$ convolutional layer for inference.
3) Fusion Block: In the fusion block, we employ the RepVGGplus block[ding2021repvgg], which leverages re-parameterization techniques to convert a multi-branch ResNet-style architecture into a single-path VGG-like model during inference. The structure of the RepVGGplus block is illustrated in Fig. 5 This maintains the representational power of the model while accelerating inference. The final deep features are aggregated as follows:
@@
where f{f} denotes the upsampling and cross-scale fusion operations.
@@ III-C Adaptive Scale-Aware 3D Regression Head
2) Scale-Semantic Feature Fusion Module: This module comprises two key components: offset feature generation and attention mask generation. For offset generation, the scale feature $G_{o}$} and local semantic features $F_{R}$} are concatenated. The combined features are processed through a stacked convolutional layer consisting of $\times3$ and $\times1$ convolutions, producing a scale-aware offset $\Delta_{S}$}, which is used to dynamically adjust the receptive field in subsequent deformable convolutions.
@@
In the attention mask generation component, inspired by[yang2021simam], we design an attention mechanism based on spatial variance. This mechanism emphasizes regions with significant deviations from the mean, which are often foreground regions of interests. First, the mean of the local semantic features $F_{R}$} are computed across spatial dimensions:
@@
3) 3D Regression Head: To achieve scale-aware dynamic receptive field adjustment, the first layer of the 3D regression head utilizes deformable convolutions[zhu2019deformable]. As shown in Eqn. 13, the offset $\Delta_{S}$} and attention mask $M$ generated by the Scale-Semantic Feature Fusion Module are applied to this layer, allowing the model to better detect and localize objects of varying scales:
@@
where $\Omega$ represents the local neighborhood for a $\times3$ kernel, $P_{n}$}} enumerates the locations in $\Omega$, $\Delta_{S}^{{P_{n}}}$} is the learned offset for each position in the convolution kernel, $W_{{P_{n}}}$}} is the weight of the convolution kernel at position $P_{n}$}}, and $M_{{P_{n}}}$}} controls the contribution of the corresponding kernel position.
Next, we normalize the feature map $F_{de}$} using the Attentive Normalization layer[li2020attentive], which is a light weight module integrating feature normalization and channel-wise feature attention:
@@
Due to AN’s mixture modeling approach to affine transformation, which recalibrates features after standardization, it is employed in the regression head to facilitate the learning of more expressive latent feature representations. Finally, the feature map is processed through a LeakyReLU activation function and a $\times1$ convolution for channel mapping, ultimately regressing the 3D bounding box attributes.